### v3

# MlingConf: A Comprehensive Study of Multilingual Confidence Estimation on Large Language Models

Abstract
The tendency of Large Language Models (LLMs) to generate hallucinations raises concerns regarding their reliability. Therefore, confidence estimations indicating the extent of trustworthiness of the generations become essential. However, current LLM confidence estimations in languages other than English remain underexplored. This paper addresses this gap by introducing a comprehensive investigation of Multilingual Confidence estimation (MlingConf) on LLMs, focusing on both language-agnostic (LA) and language-specific (LS) tasks to explore the performance and language dominance effects of multilingual confidence estimations on different tasks. The benchmark comprises four meticulously checked and human-evaluate high-quality multilingual datasets for LA tasks and one for the LS task tailored to specific social, cultural, and geographical contexts of a language. Our experiments reveal that on LA tasks English exhibits notable linguistic dominance in confidence estimations than other languages, while on LS tasks, using question-related language to prompt LLMs demonstrates better linguistic dominance in multilingual confidence estimations. The phenomena inspire a simple yet effective native-tone prompting strategy by employing language-specific prompts for LS tasks, effectively improving LLMs’ reliability and accuracy on LS tasks.
MlingConf: A Comprehensive Study of Multilingual Confidence Estimation on Large Language Models
Boyang Xuethanks:   Equal Contribution. , Hongru Wang♠∗, Rui Wang, Sheng Wang Zezhong Wang, Yiming Du, Bin Liang, Kam-Fai Wong The Chinese University of Hong Kong The University of Hong Kong {byxue, hrwang, kfwong}@se.cuhk.edu.hk
1 Introduction
LLMs’ susceptibility to generating hallucinated contents in their generations incurs concerns about unreliability in real-world applications Ji et al. (2023); Rawte et al. (2023). Therefore, it becomes increasingly crucial for users to directly ascertain how much they can trust a model’s response. Assessing the confidence or uncertainty of a model’s output can immediately indicate the level of reliability to users, thereby playing a key role in developing trustworthy AI systems Geng et al. (2023); Kadavath et al. (2022).
Figure 1: Examples of generations and confidence scores of Llama-3.1 given the same inputs in three languages in LA and LS scenarios derived from SciQ and LSQA datasets respectively.
However, existing research on confidence or uncertainty estimations of LLMs has been confined exclusively in English Kadavath et al. (2022); Lin et al. (2022); Geng et al. (2023); Tian et al. (2023b). Considering the LLMs’ global deployment, the dearth of confidence estimations in languages other than English hinders users from ascertaining the reliability of generations in non-English scenarios. Owing to the disparities in the quantity and domain coverage of pre-training corpora across different languages, the confidence estimation ability may also presumably vary. The performance of confidence estimation methods mainly employed in English remains a subject worthy of exploration when generalized to other languages.
Additionally, linguistic dominance referring to that one language holds a superior status over others within a specific social or cultural context (Blommaert, 2010; Treffers-Daller, 2019; Heller, 2007), also exist in confidence estimation ability on different languages. To achieve more reliable confidence estimations in multilingual contexts over various tasks, we study and divide the tasks into language-agnostic (LA) (Zhao et al., 2020) and language-specific (LS) scenarios as in Figure 1. In this study, the LS refers to the tasks that hold linguistic dominance caused by the knowledge domain coverage varying in different language training corpora, such as questions pertaining to social, cultural, or geographical contexts for a specific language, while the LA task involves linguistic dominance mainly caused by the quantities of training corpora, such as general knowledge, common sense, and reasoning (Basaj et al., 2018; Sánchez et al., 2024).
To this end, we propose a benchmark called MlingConf (Multilingual Confidence) to investigate the performance of several LLM confidence estimation methods on five languages including English, Japanese, Chinese, French, and Thai on both LA and LS tasks respectively. First, we meticulously construct the MlingConf dataset for this study including five high-quality multilingual datasets in LA and LS scenarios. The LA involves four tasks including closed-book factual knowledge question-answering (QA) derived from TriviaQA Joshi et al. (2017), math reasoning from GSM8K (Cobbe et al., 2021), common sense reasoning from CommonsenseQA (Talmor et al., 2019), and professional scientific knowledge QAs from (Johannes Welbl, 2017). These four datasets widely used in confidence estimation in English (Kuhn et al., 2023; Xiong et al., 2024) are translated into other four languages. We also create a language-specific QA (LSQA) for the LS scenario, which is comprised of specific social, cultural, and geographical QA pairs for the investigated five languages respectively. To ensure the data quality, we conduct rigorous translation consistency checks to filter the failed samples and finally employ linguistic experts for human evaluation.
Experiments are conducted on three major LLM confidence estimation methods including probability-based (Vazhentsev et al., 2023; Varshney et al., 2023), $p(\mathrm{True})$-based (Kadavath et al., 2022), and self-verbalized (Lin et al., 2022; Xiong et al., 2024) confidence estimations using the curated five multilingual datasets on several LLMs. We evaluate the confidence estimation ability and calibration using AUROC and ECE. Results on LA tasks suggest that all the confidence estimation methods demonstrate comparable performance of reliability across different languages, while the calibration performance in English significantly outperforms others. Results on the LS task reveal a pronounced phenomenon of language dominance, indicating that, for questions related to specific linguistic contexts, utilizing the respective languages yields the highest accuracy and confidence estimation performance. This observation inspires a native-tone prompting strategy, whereby, in the LS task, the relevant linguistic background of the question is first assessed, and then the corresponding language is employed for the response. Compared to the use of any single language, this approach leads to significant improvements in both accuracy and confidence estimation.
The contributions are summarized as follows:
$\bullet$ To the best of our knowledge, the MlingConf first propose to investigate multilingual confidence estimations with intricately constructed and expert-checked MlingConf datasets for both LA and LS scenarios, serving as a valuable benchmark for future works  [^1] The codes and the MlingConf datasets have been released on https://github.com/AmourWaltz/MlingConf.
$\bullet$ We conduct experiments on MlingConf datasets and the results reveal that language dominance affects confidence estimation performance in English on LA tasks and query-related languages on LS tasks respectively.
$\bullet$ We propose a native-tone prompting strategy, which significantly enhances the reliability and accuracy compared to the use of any single language prompts for LS tasks.
2 MlingConf Dataset
2.1 Data Source
Owing to the lack of multilingual resources to comprehensively exhibit confidence estimation across diverse languages, our study leverages several datasets in English (en) and translates into four languages, encompassing Japanese (ja), Chinese (zh), French (fr), and Thai (th). We evaluate the quality of confidence estimates on our proposed MlingConf dataset including four datasets for the language-agnostic (LA) scenario and one dataset for the language-specific (LS) scenario.
Language-Agnostic (LA) Tasks
For LA tasks, we employ the following four datasets that are widely used for confidence estimation research in English in Kuhn et al. (2023); Xiong et al. (2024). We translate the datasets into other four languages. The datasets include 1) TriviaQA (TVQA) (Joshi et al., 2017) of closed-book trivia question-answering pairs to gauge models’ factual knowledge; 2) GSM8K (Cobbe et al., 2021) for arithmetic reasoning task of math problems; 3) CommonsenseQA (CSQA) (Talmor et al., 2019) of multiple-choice QA pairs requiring different types of commonsense knowledge; 4) SciQ (Johannes Welbl, 2017) requiring scientific professional knowledge.
Language-Specific (LS) Tasks
We create Language-Specific QA (LSQA) dataset pertaining to language-dominant knowledge covering specific social, geographical, and cultural language contexts for the UK & US, France, China, Japan, and Thailand respectively. We prompt GPT-4 OpenAI (2023)  [^2] https://platform.openai.com/docs/api-reference to generate 200 questions pertaining to only one specific language contexts as a language-specific subset. As demonstrated in Figure 12, for example, all questions in Japanese subset pertain to Japanese social culture, history, or geography. All the subsets are also translated into other four languages.
The language selection, dataset details, prompts, and samples extracted from the LSQA are demonstrated in Appendix A, B, and C respectively.
Figure 2: Three stages of MlingConf dataset construction.
2.2 Dataset Construction
The construction of the MlingConf dataset in this study follows a three-stage procedure as delineated in Figure 2.
Stage 1
: The QA samples derived from the above datasets are translated into four languages (ja, th, zh, fr) through GPT-4 OpenAI (2023);
Stage 2
: We check the consistency of five translated results by comparing the semantic equivalence in pairs in $C_{5}^{2}=10$ times for each sample. The samples with more than 2 times semantic in-equivalence are treated as noisy data and then filtered. The changes of samples before and after consistency check and filter are shown in Table 1 and more clean multilingual datasets are obtained. Moreover, we present the number of samples for each language-specific subset in the LSQA dataset as in Table 2.
TVQA GSM8K CSQA SciQ LSQA
Before 2000 1319 1221 1000 1000
After 1238 1318 1152 640 857
Table 1: Number of samples before and after consistency check and filter.
LSQA en zh ja fr th Total
185 172 167 179 154 857
Table 2: Statistics of samples for each language-specific subset of the LSQA dataset.
Stage 3
: Finally, we employ several experts majoring in linguistics to examine the translation performance across 50 randomly selected samples as shown in Table 3. Human evaluation results suggest the obtained multilingual datasets are high-quality for further experiments. Temperature $T$ is set to 0 for all the generations. The translation and semantic equivalence comparison prompts are presented in Appendix C.
Lang. TVQA GSM8K CSQA SciQ LSQA
zh 96% 100% 100% 98% 100%
ja 98% 100% 98% 96% 100%
fr 100% 100% 100% 98% 100%
th 96% 100% 98% 94% 100%
Table 3: Translation accuracy evaluated by linguistic experts on 50 randomly selected samples.
3 Experimental Settings
3.1 Confidence Estimation Methods
In this part, we investigate three confidence estimation methods primarily used in English for LLMs as in Figure 6. These methods will be also conducted in our multilingual settings. Specifically, we denote $\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})$ as the confidence score associated with the output sequence $\boldsymbol{y}=[y_{1},y_{2},\dots,y_{N}]$ given the input context $\boldsymbol{x}=[x_{1},x_{2},\dots,x_{M}]$.
Probability-based Confidence (Prob.):
The probability-based confidence is estimated by calculating the joint token-level probabilities over $\boldsymbol{y}$ conditioned on $\boldsymbol{x}$. As longer sequences are supposed to have lower joint likelihood probabilities that shrink exponentially with length, we calculate the geometric mean by normalizing the output token probabilities which are represented as:
$\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})$ $=\left({\prod^{N}_{i}p(y_{i}|\boldsymbol{y}_{<i},\boldsymbol{x})}\right)^{\frac{1}{N}}$ (1)
$p(\mathrm{True})$-based Confidence (p(True)):
The $p(\mathrm{True})$ confidence score is implemented by simply asking the model itself if its first proposed answer $\boldsymbol{y}$ to the question $\boldsymbol{x}$ is true Kadavath et al. (2022), and then obtaining the probability $p(\mathrm{True})$, which can implicitly reflect self-reflected certainty.
Self-verbalized Confidence (Verb.):
As LLMs possess good self-reflection and instruction-following abilities, recent works pay particular attention to linguistic confidence via prompting LLMs to express certainty in verbalized numbers or words Lin et al. (2022); Xiong et al. (2024). We adopt verbalized numerical probability in token-level space as LLM’s confidence estimation.
The multilingual prompts for p(True) and Verb. are in Appendix C.
3.2 Evaluation Metrics
Accuracy (Accu.)
We employ a string-matching approach to evaluate the accuracy of generated answers $\boldsymbol{y}$ and compare them with the ground truth $\boldsymbol{\hat{y}}$. Although exact matching (EM) of $\boldsymbol{y}\equiv\boldsymbol{\hat{y}}$ is widely used on GSM8K and CSQA, it always misjudges some correct answers with slight differences on closed-book QA tasks, to better assess the result accuracy (Accu.), we replace EM with a variant called positive-recall exact matching (PREM) of $\boldsymbol{y}\in\boldsymbol{\hat{y}}\vee\boldsymbol{\hat{y}}\in\boldsymbol{y}$. Comparisons of several EM variants we tested as well as human evaluations are presented in Appendix D.
Area Under the Receiver Operator Characteristic Curve (AUROC)
AUROC assesses the effectiveness of confidence estimation Filos et al. (2019) by quantifying how likely a randomly chosen correct answer possesses a higher confidence score than an incorrect one, yielding a score within the range of [0, 1], implemented by sklearn toolkit  [^3] https://github.com/scikit-learn/scikit-learn/blob/main/sklearn/metrics/_ranking.py. A higher AUROC score implying higher reliability is preferred.
Expected Calibration Error (ECE)
ECE gauges the calibration performance which indicates how well a model’s predicted confidence matches its actual accuracy (Guo et al., 2017a). For an expected well-calibrated AI system, samples $\boldsymbol{x}$ with confidence of $q$ should also have an average accuracy of $q$ on predictions $\boldsymbol{y}$ where $P(\boldsymbol{y}=\boldsymbol{\hat{y}}|\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})=q)=q$ with ECE=0. ECE is essential for reliable AI systems on prediction tasks like weather forecasting. The smaller the ECE value, the better. Details of the ECE calculation are presented in Appendix D.
3.3 Implementation Details
The experiments are conducted on three LLMs: GPT-3.5-Turbo-0125 (GPT-3.5), Llama-3.1-8B-Instruct (Llama-3.1)  [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct (Dubey et al., 2024), and Llama-2-13B-Chat (Llama-2)  [^5] https://huggingface.co/meta-llama/Llama-2-13b-chat (Touvron et al., 2023), and Vicuna-7B-v1.5 (Vicuna-v1.5)  [^6] https://huggingface.co/lmsys/vicuna-7b-v1.5 (Zheng et al., 2023). We only present the results of the current most commonly used commercial LLM GPT-3.5 and open-source LLM Llama-3.1 in the main part and leave others in Appendix E. Few-shot prompts containing $N_{f}$ examples are utilized for answer generation with the greedy decoding where $N_{f}$ is set to 8 for GSM8K and 4 for others.
4 Experimental Results and Analysis
Dataset Conf. en zh ja fr th Avg.
ARC. $\uparrow$ ECE $\downarrow$ ARC. $\uparrow$ ECE $\downarrow$ ARC. $\uparrow$ ECE $\downarrow$ ARC. $\uparrow$ ECE $\downarrow$ ARC. $\uparrow$ ECE $\downarrow$ ARC. $\uparrow$ ECE $\downarrow$
GPT-3.5
TVQA Prob. 76.51 24.36 78.39 32.95 76.90 28.14 72.14 27.39 74.30 40.17 75.65 30.60
p(True) 79.64 18.25 82.34 22.94 84.50 29.06 80.59 20.90 81.22 40.87 81.66 26.40
Verb. 80.32 16.52 81.76 24.32 84.61 34.49 83.47 26.53 86.72 38.19 83.38 28.01
GSM8K Prob. 54.79 26.48 58.49 27.19 57.09 29.46 57.38 37.29 61.73 41.77 57.90 32.44
p(True) 65.25 31.88 62.74 28.65 69.75 19.03 60.14 39.08 61.45 49.88 63.87 33.70
Verb. 62.34 22.17 59.25 28.91 58.34 26.71 66.65 25.14 54.02 45.63 60.12 29.71
CSQA Prob. 59.06 24.45 55.92 38.30 48.01 50.60 55.33 31.12 48.21 41.71 53.31 48.18
p(True) 67.13 19.65 58.64 27.08 65.23 19.24 66.33 23.43 59.96 34.47 63.46 24.77
Verb. 69.60 16.84 54.30 21.54 68.34 19.84 61.87 21.81 68.93 21.71 64.73 24.35
SciQ Prob. 69.50 32.23 71.29 35.63 78.28 47.06 72.66 34.85 75.13 56.17 73.37 41.19
p(True) 72.06 23.15 76.18 30.44 80.16 36.18 71.30 37.85 68.29 41.25 73.60 33.77
Verb. 70.18 20.80 75.50 37.59 77.89 30.33 69.31 32.47 74.85 41.15 73.55 32.47
Avg. Prob. 64.96 26.88 66.02 33.51 65.07 38.71 64.37 32.66 64.88 44.95 65.06 35.34
p(True) 70.38 23.23 69.97 27.27 74.91 25.87 69.59 30.31 67.73 41.61 70.65 29.66
Verb. 70.61 19.01 67.70 28.09 72.29 27.84 70.32 26.48 71.13 36.67 70.45 28.64
Overall 68.68 23.04 67.90 29.65 70.75 30.84 68.10 29.82 67.90 41.08 68.66 30.89
Llama-3.1
TVQA Prob. 80.74 10.72 80.41 40.39 88.75 26.24 79.05 20.38 89.59 36.77 83.71 26.90
p(True) 68.98 18.35 68.10 38.19 52.69 37.85 62.00 22.04 60.55 37.01 62.66 30.69
Verb. 77.18 24.73 63.50 37.64 68.91 34.27 69.90 25.44 73.22 40.19 70.54 32.45
GSM8K Prob. 65.69 22.33 66.37 21.92 69.73 35.69 61.07 29.56 63.22 28.51 65.22 27.60
p(True) 61.64 14.49 65.83 17.39 71.40 11.26 57.04 8.02 57.31 12.90 62.64 12.81
Verb. 57.00 50.05 63.04 42.89 58.93 45.33 54.45 35.31 55.30 34.94 57.75 41.70
CSQA Prob. 78.06 13.64 64.91 36.33 75.65 30.11 66.12 19.72 77.65 42.18 72.47 28.40
p(True) 56.25 34.04 64.24 36.82 66.60 40.32 59.34 27.72 58.07 29.91 60.90 33.76
Verb. 62.42 28.16 54.61 28.84 61.06 37.85 57.97 23.96 71.91 37.15 61.39 31.19
SciQ Prob. 74.14 13.40 72.09 32.26 74.48 34.21 77.45 22.76 77.61 36.10 75.15 27.75
p(True) 62.38 19.28 64.89 37.01 58.92 36.47 61.01 10.72 51.90 41.06 59.82 28.91
Verb. 62.65 24.10 52.90 32.94 69.10 39.15 59.30 24.92 65.93 40.67 61.98 38.36
Avg. Prob. 74.88 15.02 70.94 32.72 77.15 31.56 70.92 23.10 77.01 35.89 74.14 27.66
p(True) 62.31 21.54 65.76 32.21 62.40 31.47 59.59 17.12 56.95 30.22 61.51 26.54
Verb. 64.81 31.76 58.51 35.57 64.50 39.15 60.40 27.40 66.59 38.23 62.92 35.93
Overall 67.26 22.77 65.10 33.55 67.97 34.06 63.72 22.58 66.85 34.78 66.19 29.38
Table 4: Experimental results of AUROC (ARC.) and ECE of three confidence estimation methods on four LA datasets on GPT-3.5 and Llama-3.1.
In the experiment part, we aim to address the following three research questions:
a. How do performance (ARC. and ECE) vary in three confidence estimation methods on LA tasks?
b. How do performance and accuracy vary in different languages on LA tasks?
c. How do performance and accuracy vary in language usage on different language-specific LSQA subsets on LS tasks?
4.1 Results and Analysis on LA Tasks
We present the results of the confidence estimation performance of four LA datasets using GPT-3.5 and Llama-3.1 as presented in Table 4. The key findings and analysis for research questions a and b are listed below.
Figure 3: Experimental results of Accuracy on four LA datasets on GPT-3.5 and Llama-3.1.
Recalling research question a, we highlight the supreme performance using   among the three methods for each column of the dataset in Table 4. On GPT-3.5, both p(True) and Verb. yield the superior performance than Prob. across all languages averaged on four datasets (ARC.: 70.65, 70.45 vs. 66.02; ECE: 29.66, 28.64 vs. 35.34). p(True) and Verb. have comparable performance in ARC. scores, while Verb. is better calibrated. In contrast, Prob. show the superior performance than p(True) and Verb. and performs more stable on Llama-3.1 (ARC.: 74.14 vs. 61.51 and 62.92; ECE: 27.66 vs. 26.54 and 35.93). p(True) demonstrates better calibration results on languages other than English.
The performance differences between GPT-3.5 and Llama-3.1 can be attributed to that GPT-3.5’s strong instruction-following abilities benefits the prompt-based confidence estimation methods Verb. and p(True)., while Llama-3.1’s likelihood probabilities are well-calibrated during training.
Recalling research question b, previous studies only validate the efficacy of prompt-based confidence assessments exclusively within English. Our findings indicate that the approaches are also preferable in other languages and the performance fluctuates in different languages. Specifically as demonstrated in Table 4, ARC. scores are less fluctuating across different languages while ECE in English (23.04 and 22.77) performs better than in other languages on both GPT-3.5 and Llama-3.1.
Despite the powerful multilingual capacity of LLMs, discrepancies exist in the quantity of distinct linguistic training corpora available for each language. The results suggest that ARC. is a metric not significantly related to language usage in LLMs, while the strong performance of ECE in English can be attributed to the extensive training corpus or the specific calibration conducted during training in English. As the only middle-resource language, Thai exhibits a notably lower level of reliability compared to the other high-resource languages.
We also report the response accuracy on LA datasets in Figure 3, English consistently performs better on the datasets exclusively GSM8K. Considering consistency check Table 1, the lowest filter rate in GSM8K translation indicates that mathematical reasoning tasks are minimally affected by language bias. As a result, the performance fluctuations across different languages on GSM8K are relatively small. For Chinese exhibit slightly superior mathematical capabilities compared to English on GSM8K on both GPT-3.5 and Llama-3.1 (Accu. 48.09 and 63.96), it is hypothesized that pre-training corpora contain a substantial amount of Chinese mathematical problems.
Additionally, GPT-3.5 and Llama-3.1 exhibit notable superiority on multilingual tasks over Llama-2 and Vicuna-1.5. Therefore, we only present the results on GPT-3.5 and Llama-3.1 in Table 4 and leave the other two in Appendix E.
Figure 4: Experimental results of AUROC and ECE of three confidence estimation methods on five language-specific subset of LSQA using Llama-3.1.
Figure 5: Experimental results of Accuracy on five language-specific LSQA subsets using Llama-3.1.
4.2 Results and Analysis on LS Tasks
We present the results of the confidence estimation performance on five language-specific subsets of the LSQA dataset using Llama-3.1 as presented in Figure 4 and 5. The key findings and analysis for research question c are listed below.
Recalling research question c, in comparison with the LA task in Sec. 4.1 where linguistic dominance is primarily manifested in English, on LSQA, linguistic dominance is determined by the language of the subset. As illustrated in Figure 4, the diagonal values of the ECE and ARC. heatmaps of Prob. are more pronounced, indicating that when using Prob. confidence, linguistic dominance is more apparent compared to p(True) and Verb.. Consequently, we have opted Prob. for subsequent experiments in 5. Additionally, as shown in Figure 5, although prompting in English performs well and stable across different subsets, there is a noticeable improvement in accuracy when prompts are related to each subset’s language.
The linguistic dominance on LSQA can be conjectured to stem from the fact that such data pertaining to the language-specific cultural, geographical, or social contexts are already included in the pre-training corpora of their respective languages, thereby achieving optimal performance when prompting their specific languages. The use of maximum likelihood estimation during LLM pre-training further accounts for Prob. confidence exhibits more prominent linguistic dominance than p(True) and Verb. confidence estimations.
5 Native-Tone Prompting Method
Confidence estimation performance differences caused by linguistic dominance on LA and LS tasks in Sec. 4 motivate us to further propose a research question How can we better leverage linguistic dominance to improve the LLMs’ reliability and accuracy on LA and LS tasks?
Method
In Sec. 4.1, prompting in English generally outperforms others, hence responding in English on LA tasks can be adequately credible and accurate where linguistic dominance is leading in English. Inspired by the linguistic dominance phenomena on each language-specific LSQA dataset in Sec. 4.2, we propose a simple yet effective Native-Tone Prompting (NTP) strategy to achieve better confidence estimation performance and accuracy on the LS tasks. NTP first prompts LLMs to identify the language context of the question, and then uses that language to answer the question, effectively exhibiting a "native tone" that is more familiar in that language-related context. The prompt of NTP is presented in Appendix C.
Prompt en zh ja fr th NTP
Accu. $\uparrow$ 77.68 60.16 44.64 60.90 35.18 79.46
ARC. $\uparrow$ 73.07 72.44 76.87 71.92 74.60 77.25
ECE $\downarrow$ 12.15 30.93 32.31 20.50 40.69 10.28
Table 5: Experimental results of overall Accu., ARC., and ECE on the LSQA dataset by prompting using different languages and our proposed NTP method.
Experiments
In Table 5, we present the results of prompting by any single language versus NTP. The experiments demonstrate that our proposed NTP better leverages the inherent linguistic dominance, thereby yielding more reliable and accurate results than any single language prompt, validating the effectiveness of NTP on LS tasks.
Discussion
Linguistic dominance exerts a significant influence on multilingual confidence estimations on both LA and LS scenarios, where the dominance is manifested in English and query-related language respectively. This phenomenon indicates that the multilingual capabilities and reliability of LLMs are still constrained by the imbalanced training corpus among diverse languages. Furthermore, the reliability and accuracy on English, serving as the primary training corpora, have not been adequately generalized to other languages. Even for semantically equivalent queries, the reliability of responses cannot be consistently maintained, highlighting challenges in enhancing multilingual reliability for LLMs.
6 Related Works
Confidence Estimation for LLMs
Both confidence and uncertainty estimations indicate the level of certainty or assurance of a response $\boldsymbol{y}$ generated by LLMs given a query $\boldsymbol{x}$ and are occasionally regarded interchangeably Geng et al. (2023). Previous confidence estimation methods can be categorized into five distinct classes, as illustrated in Figure 6. ➀ Probability-based method: Following model calibration on classification tasks (Guo et al., 2017b), Vazhentsev et al. (2023) intermediately quantify sentence uncertainty over token probabilities; ➁ $p(\mathrm{True})$-based method: Kadavath et al. (2022) instructs the LLM to self-evaluate the correctness of the generated answer by directly accessing $p(\mathrm{True})$; However, both ➀ and ➁ require access to token probabilities and thus are limited to white-box LLMs. ➂ Self-verbalized method: Recently, LLMs’ remarkable instruction-following ability provides a view of expressing confidence in words (Lin et al., 2022; Zhou et al., 2023; Tian et al., 2023a; Xiong et al., 2024) especially for black-box LLMs; ➃ Sampling-based method: By sampling multiple responses to a given question, Xiong et al. (2024); Lyu et al. (2024) aggregate all the confidence scores as the indicator. Kuhn et al. (2023) proposes semantic entropy to quantify uncertainty for sequences with shared meanings at the semantic level. ➄ Training-based method: Mielke et al. (2022); Lin et al. (2022); Han et al. (2024); Kadavath et al. (2022) propose to train an external natural language inference (NLI) module or LLM itself to improve confidence estimations.
Current research for LLM confidence estimations is focused on English. We first present a comprehensive study to investigate LLMs’ confidence estimations in languages other than English. Note that sampling- and training-based methods are omitted in this work as they are cost-expensive and time-consuming for multilingual tasks.
Multlingual LLMs
Most recent LLMs primarily pre-trained on English corpora have showcased remarkable capabilities Pires et al. (2019); Shi et al. (2023); OpenAI (2023); Workshop et al. (2023); Chowdhery et al. (2022); Zhang et al. (2022); Brown et al. (2020). However, their efficacy in other low-resource languages remains limited, hindering the global proliferation of LLM deployment and provoking a surge in research interest. Recently, many research works have extended various tasks in multilingual domains such as claim fact-checking (Pikuliak et al., 2023) and jailbreak problem (Deng et al., 2024). Prior studies have also underscored the advantages of leveraging pre-trained multilingual models in diverse cross-lingual applications Wang et al. (2023a, b); Bhattacharjee et al. (2023); Qin et al. (2022). Inspired by them, our research systematically delves into an under-explored domain of multilingual confidence estimation.
7 Conclusion
This study underscores the necessity of advancing multilingual confidence estimation methods for LLMs to ensure their reliability across diverse linguistic contexts. The proposed MlingConf serves as a valuable and noteworthy benchmark to address the gap in multilingual confidence estimation research. Our findings demonstrate the variability of performance of different confidence estimation methods on both LA and LS scenarios, revealing the influence of linguistic dominance on different tasks. This leads to the NTP strategy, improving accuracy and reliability by aligning the response language with the linguistic context of the query for LS tasks. These insights and the introduction of MlingConf datasets pave the way for future research to refine these methods, enhancing the global applicability and trustworthiness of LLMs.
Limitations and Future Work
The limitations and prospects for future research are outlined as follows:
Extensions on More Low-Resource Languages
The present study is constrained by the substantial cost associated with the API cost using GPT-4 for translation as well as linguistic verification. This multilingual research is restricted to five languages in the first version. This initial phase aims to delve into confidence estimation within multilingual domains. Our future endeavors will involve the expansion of the benchmark dataset, encompassing additional languages and data points to enrich our investigations.
Extensions on Native-Tone Prompting
Although the proposed Native-Tone Prompting method can enhance the accuracy and reliability of LS tasks, it still relies on external prompts to determine which language domain the query pertains to. Moving forward, it is promising to broaden the scope of developing a cross-lingual method that can directly transfer the specific language dominance to other language contexts, thereby facilitating multilingual confidence estimation abilities for LLMs.
Ethics Statement
In this paper, we introduce several self-constructed multilingual datasets derived from the publicly available dataset. The selection of investigated languages in this work depends on whether we can employ appropriate linguistic experts. Most linguistic specialists are M.Phil. or Ph.D. students majoring in linguistics and others are from crowd-sourcing platforms. We meticulously adhered to legal and ethical standards throughout the data collection process, prioritizing privacy and obtaining informed consent. Linguistic experts were furnished with comprehensive details regarding the study’s objectives, data collection methodologies, and associated risks or benefits. They were afforded the opportunity to seek clarifications and voluntarily provide consent before their involvement. All collected data were exclusively utilized for research purposes.
Appendix A Language Information
The basic information of ISO codes and the language family of the investigated languages is presented in Table 6. The investigated languages from widely spoken to lesser-known ones in this work are selected following three principles.
1) Following (Lai et al., 2023; Deng et al., 2024) which determines the resource levels for each language by utilizing the data ratio from the CommonCrawl corpus  [^7] http://commoncrawl.org/, we select three languages (Chinese, Japanese, and French) in high-resource category whose data ratio exceeds 1%, and one language (Thai) from medium-resource class that falls between 0.1% and 1%. To ensure the confidence estimation ability can be observed, the low-resource languages less than 0.1% are omitted and left for future works.
2) This selection ensures coverage of a wide range of linguistic characteristics from different language families as in Table 6. A language family represents a collective of cognate languages stemming from a common ancestral source, serving as a focal point within the domain of linguistics  [^8] https://en.wikipedia.org/wiki/Language_family.
3) For each selected language, we can employ one linguistic expert for the human check to ensure the data quality;
ISO 639-1 Family
English en Indo-European
French fr Indo-European
Chinese zh Sino-Tibetan
Japenese ja Japanese-Ryukyuan
Thai th Kra–Dai
Table 6: List of International Standard Organization (ISO) 639-1 codes and language family information.
Appendix B Dataset Details
TriviaQA
The TriviaQA dataset Joshi et al. (2017) is a realistic text-based reading comprehension question-answering dataset containing 650K question-answer-evidence triples from 95K documents collected from Wikipedia and the websites, served as a benchmark for evaluating machine comprehension and question-answering systems, which is more challenging than standard QA benchmark datasets where the answer spans can be directly retrieved and copied.
GSM8K
GSM8K (Grade School Math 8K) (Cobbe et al., 2021) is a dataset of 8.5K high quality linguistically diverse grade school math word problems. The dataset was created to support the task of question answering on basic mathematical problems that require multi-step reasoning to solve.
CommonsenseQA
CommonsenseQA (Talmor et al., 2019) is a new multiple-choice question answering dataset that requires different types of commonsense knowledge to predict the correct answers. The dataset consists of 12,247 questions with 5 choices each.
SciQ
The SciQ dataset (Johannes Welbl, 2017) contains 13,679 crowdsourced science exam questions about Physics, Chemistry and Biology, among others. The questions are in multiple-choice format with 4 answer options each. For the majority of the questions, an additional paragraph with supporting evidence for the correct answer is provided.
LSQA
We present two examples of the LSQA dataset in English- and Japanese- specific subsets in Figure 12.
Appendix C Prompt Details
The translation prompt for multilingual dataset construction and semantic equivalence comparison prompt for consistency check in Sec, 2 are presented in 7 and 8 respectively. Standard multilingual Question-Answering prompts are in 9. Multilingual confidence estimations of P(True) and Verb. are presented in Fig. 10 and 11. Notably, the prompts for self-reflected true probability confidence estimation are followed by previous work Kadavath et al. (2022); Kuhn et al. (2023).
Figure 7: Translation prompt.
Figure 8: Semantic equivalence comparison prompt.
Figure 9: Multilingual Question-Answering prompts.
Figure 10: Multilingual p(True)-based Confidence Estimation Prompt.
Figure 11: Self-verbalized confidence estimation prompt.
Figure 12: Examples of the LSQA ataset in English and Japanese specific subsets.
Figure 13: Native-tone prompting (NTP).
Appendix D Metric Details
Expected Calibration Error (ECE)
We partition the inference results into $M$ disjoint bins $\left\{B_{m}\right\}_{m=1}^{M}$ based on the confidence scores $\left\{q_{i}\right\}$, compute the average confidence score in $\left(\frac{m-1}{M},\frac{m}{M}\right]$ for the $m$-th bin $B_{m}$, and compare it with the average true accuracy $\mathsf{acc}(B_{m})$ of the answers within $B_{m}$. The ECE is calculated by:
$\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\mathsf{acc}(B_{m})-\mathsf{conf}(B_{m})\right|$ (2)
The average accuracy $\mathsf{acc}(B_{m})$ and confidence $\mathsf{conf}(B_{m})$ of the answers in $B_{m}$ is obtained by:
$\mathsf{acc}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{a_{i}\in_{m}}\mathbb{I}(\hat{b}_{i}=b_{i})$ (3)
$\mathsf{conf}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{a_{i}\in_{m}}q_{i}$ (4)
where $a_{i}$, $b_{i}$, $\hat{b}_{i}$, and $q_{i}$ indicate the input data, label, prediction result, and confidence score respectively for the $i$-th sample.
Accuracy
For closed-book QA evaluation, we observe that simply applying EM may misjudge the correct answers. We compare several variants of EM as in Table 7 and report their successful judgments on responses of 20 selected samples that are misjudged using EM, where PEM, RRM, and PREM indicate Positive-EM, Recall-EM, and Positive-Recall-EM and the mathematical explanations are presented in Table 7. Upon human discrimination, EMPR exhibits the lowest failure rate and is therefore selected as the evaluation metric for this work.
Variant Explanation # Fail
EM $\boldsymbol{y}\equiv\boldsymbol{\hat{y}}$ 20
PEM $\boldsymbol{y}\in\boldsymbol{\hat{y}}$ 16
REM $\boldsymbol{\hat{y}}\in\boldsymbol{y}$. 6
PREM $\boldsymbol{y}\in\boldsymbol{\hat{y}}\vee\boldsymbol{\hat{y}}\in\boldsymbol{y}$. 2
Table 7: Number of failed judgments by human check for different EM variants.
Appendix E Appendix Experiments
We present the experimental results on Llama-2 and Vinuca-1.5 in Table 8. Results suggest that confidence estimation abilities are relatively weak on both Llama-2 and Vinuca-1.5 across three methods.
Dataset Conf. en zh ja fr th
ARC. $\uparrow$ ECE $\downarrow$ ARC. $\uparrow$ ECE $\downarrow$ ARC. $\uparrow$ ECE $\downarrow$ ARC. $\uparrow$ ECE $\downarrow$ ARC. $\uparrow$ ECE $\downarrow$
Llama-2
TVQA Prob. 51.92 20.56 51.24 34.12 51.92 32.04 49.51 29.33 49.63 48.72
p(True) 55.89 17.09 82.65 46.11 81.76 43.47 65.79 29.59 70.62 55.99
Verb. 59.78 21.10 53.20 45.71 51.95 39.33 61.66 38.98 54.23 59.80
GSM8K Prob. 42.72 32.18 50.46 33.59 50.35 51.67 43.60 36.25 55.11 52.87
p(True) 60.82 49.30 59.86 58.87 62.89 67.39 59.51 62.36 47.91 75.22
Verb. 59.39 43.65 53.29 54.59 53.40 49.27 53.26 37.61 54.53 56.96
CSQA Prob. 49.30 30.40 49.95 31.72 50.28 43.28 49.72 27.40 50.23 40.84
p(True) 56.53 26.05 55.34 45.65 53.46 46.01 59.76 25.49 50.21 63.09
Verb. 53.64 19.54 51.74 24.06 50.36 34.03 52.93 15.08 50.73 62.01
SciQ Prob. 55.40 24.65 76.39 44.42 74.97 45.56 62.32 39.76 51.93 59.05
p(True) 48.60 32.18 52.02 40.44 51.60 30.19 49.53 32.50 45.26 43.75
Verb. 56.34 19.89 55.17 41.36 55.58 37.20 60.27 39.14 71.17 54.95
Vicuna-1.5
TVQA Prob. 45.45 35.34 48.43 47.07 51.75 36.63 46.13 35.19 53.17 40.73
p(True) 47.45 23.58 78.96 42.86 79.71 42.40 60.38 28.89 76.58 53.43
Verb. 55.74 21.41 52.98 57.36 50.94 54.89 55.46 42.76 45.86 71.83
GSM8K Prob. 50.90 53.91 51.07 49.00 50.51 53.73 50.42 49.08 50.19 55.40
p(True) 65.40 68.30 67.28 59.33 51.09 60.70 66.78 55.69 52.86 60.83
Verb. 55.66 46.26 53.90 48.75 54.60 48.03 53.70 45.62 61.81 51.87
CSQA Prob. 48.88 26.04 50.01 43.65 49.67 45.64 45.94 31.53 49.78 51.78
p(True) 65.00 27.06 57.39 35.80 57.62 38.54 48.21 25.95 50.53 55.54
Verb. 52.32 29.80 52.29 38.59 51.08 44.49 58.68 35.15 51.90 61.77
SciQ Prob. 38.10 50.94 48.69 44.44 50.07 42.19 38.65 37.26 49.75 49.70
p(True) 45.78 31.29 73.55 40.17 66.85 45.15 59.20 42.18 74.16 58.34
Verb. 55.13 36.47 51.66 55.27 51.92 56.98 56.33 57.93 52.89 57.74
Table 8: Experimental results of AUROC (ARC.) and ECE of three confidence estimation methods on four LA datasets on GPT-3.5 and Llama-3.1.